### v1

@@ Abstract
Human walkers traverse diverse environment and demonstrate different gait locomotion and energy cost on granular terrains compared to solid ground. We present a stiffness-based model predictive control approach of knee exoskeleton assistance on sand. The gait and locomotion comparison is first discussed for human walkers on sand and solid ground. A machine learning-based estimation scheme is then presented to predict the ground reaction forces (GRFs) for human walkers on different terrains in real time. Built on the estimated GRFs and human joint torques, a knee exoskeleton controller is designed to provide assistive torque through a model predictive stiffness control scheme. We conduct indoor and outdoor experiments to validate the modeling and control design and their performance. The experiments demonstrate the major muscle activation and metabolic reductions by respectively $15$% and $3.7$% under the assistive exoskeleton control of human walking on sand.
Index Terms:
@@
Human locomotion, granular terrain, exoskeleton control, model predictive control
I Introduction
@@
Humans often traverse real-world environments with a variety of terrains that can disrupt steady gait and require additional effort [1]. While walking on solid ground is well-studied, navigating granular terrains like sand poses significant challenges [2]. Studies have shown that walking on sand requires large ankle plantarflexion in stance phase and increased knee and hip flexion during leg swing [3]. Substantial changes on walking kinematics and greater postural instability were observed for walking on sand [4]. Similar studies of walking on compliant substrates demonstrated elevated activity and great mechanical work through large excursion and maximum flexion in hip and knee joints [5]. Recent work in [6] offered ground reaction forces and gait datasets with biomechanics of walking on solid ground and sand.
@@
Although walking on granular terrain (e.g., sand) has been explored, limited studies were reported for using wearable assistive devices such as exoskeletons for enhancing the walking stability and reducing metabolic expenditure. Recent advancements in wearable exoskeletons and prosthesis development have shown promising capability in restoring ambulatory function [7], aiding postural adjustment, reducing muscle load [8], and decreasing metabolic cost during locomotion [9]. Various control methods have been developed, including quasi-stiffness approaches that match target profiles [10]. Recent advancements include reinforcement learning for impedance control parameter configuration in knee prostheses [11]. Novel impedance modulation strategies have been proposed for lower-limb exoskeletons to assist sit-to-stand movements while preserving wearer control [12], and for soft exoskeletons to aid push-off on various terrains [13]. Trajectory tracking controls have been applied to lower limb exoskeletons to personalize and stabilize user gait movements with adaptive strategies [14, 15]. Model predictive control (MPC) has been developed for compliant device behavior [16] and on-the-fly transitions between assistance modes [17, 18].
@@
All aforementioned exoskeleton applications were mainly for human walking on solid ground and the developed controllers cannot be directly applied to granular terrains. To address the knowledge gap, we focus on designing and evaluating a knee exoskeleton for walking assistance on sandy terrains, aiming to enhance ambulation ability while reducing metabolic cost. We propose a model predictive stiffness control design of human-exoskeletons to provide assistive knee joint torque and regulate walking gait on sand. A stiffness-based computation of desired knee torque is first designed using the observations of walking locomotion on sand and solid ground. A neural network is used to model human foot-terrain interactions and compute resultant ground reaction forces (GRFs) from measurements by wearable inertial measurement units (IMUs). Both desired knee torque and external GRFs are treated as inputs of the human-exoskeleton dynamic model. Finally, an MPC framework is formulated to regulate assistive torque throughout the gait phase. Both indoor terrain and outdoor experiments are conducted to validate the controller and demonstrate its efficacy.
The main contribution of this work is twofold. To the authors’ best knowledge, this is the first study to propose a stiffness-based MPC for knee exoskeletons assistive control that is specifically designed for walking on sand. Human walking dynamics on yielding terrains such as sand differ significantly from that on solid ground, and the proposed control strategy addresses these unique challenges. Second, the integration of a machine learning-based approach to estimate GRFs in real-time using wearable IMUs enables the exoskeleton controller to adapt to varying terrain conditions. This real-time adaptability enhances the system’s ability to regulate gait, providing effective assistance in complex environments.
II Human Walking on Granular Terrain
II-A Gait and Locomotion on Granular Terrains
Compared to solid ground, human walking on sand shows different gait and locomotion characteristics. Fig. 1 shows the human subject with wearable IMUs and bilateral knee exoskeletons. Fig. 1 illustrates the kinematics and modeling configuration of the human-exoskeleton interactions on sand.
@@ II-A Gait and Locomotion on Granular Terrains
Using experimental data from [6], we compare the biomechanical features of walking on sand versus solid ground. Fig. 2 shows the knee angle (denoted by $\theta_{k}$) and joint moment (denoted by $\tau_{h}$) over the gait phase (denoted by $s$) for walking on both terrains. Fig. 2 compares the GRFs from experiments with $20$ human subjects. The comparative analysis reveals that when walking on sand, knee moments show increased angular displacement and decreased amplitude. The longitudinal force $F_{x}$ and vertical force $F_{z}$ clearly exhibit distinct patterns when walking on sand compared to solid ground. Although the hip and ankle joints contribute to propulsion and mechanical work during walking on sand, the knee joint plays a key role in stabilizing and controlling gait, especially on granular terrain. By enhancing knee stability, the overall energy is reduced and gait efficiency is improved.
@@
II-B Learning-based GRF Estimation
We use a machine learning-based model to predict the GRFs. The model utilizes both temporal and spatial features from IMU measurements, offering a practical, non-invasive solution for real-time GRF estimation on sand by capturing motion and force data without direct ground contact. Fig. 3 illustrates a dual-pathway machine learning model for real-time GRF prediction. Input data are from IMUs mounted on the stance leg’s shank, heel, and toe; see Fig. 1. Similar to [19], only the accelerations in the sagittal plane ($x$- and $z$-axis) and the frontal-axis gyroscopic measurements ($y$-axis) from each IMU are used as the input for the machine learning model, i.e., ${\boldsymbol{X}}_{t}=[a_{x}(t)\;a_{z}(t)\;\omega_{y}(t)]^{T}$ from all three IMUs at time $t$.
@@ II-B Learning-based GRF Estimation
Outputs from LSTM and MLP are combined via dot product, emphasizing temporal-spatial feature interaction for a more accurate GRF estimation. A final dense layer with linear activation predicts GRFs. The model is trained using mean squared error loss and Adam optimizer, with early stopping and learning rate scheduling for regularization and optimization. This architecture is hypothesized to outperform traditional single-pathway models [20] by explicitly modeling both temporal progression and spatial configuration of movements.
III Stiffness-Based Exoskeleton Control
Fig. 4 illustrates the overall design for the knee exoskeleton control. A two-link limb model is used with joint angle ${\boldsymbol{q}}=[\theta_{t}\;\theta_{k}]^{T}\in\mathbb{R}^{2}$, where $\theta_{t}$ and $\theta_{k}$ represent the thigh and knee angle, respectively. The exoskeleton control is built on the real-time estimation of the human torque $\hat{{\boldsymbol{\tau}}}_{h}$ from the stiffness-based model and the external torque $\hat{{\boldsymbol{\tau}}}_{ext}$ from the machine learning model of the GRFs. We use the IMUs and encoders to obtain the gait phase variable $s$ and the lower-limb joint kinematics.
@@ III Stiffness-Based Exoskeleton Control
With the estimates $\hat{{\boldsymbol{\tau}}}_{h}$ and $\hat{{\boldsymbol{\tau}}}_{ext}$, an MPC scheme is formulated to design the controlled exoskeleton torque, denoted by ${\boldsymbol{\tau}}_{e}$. The lower-level controller of the exoskeletons takes the MPC outcome to drive the human-exoskeleton system to maintain desired joint angle and torque profiles on different terrains. The exoskeleton lower-level control is designed as a proportional-integral (PI) controller and the feedback torque signal is from the motor current sensor.
@@
IV Experimental Design
IV-A Experiment Setup and Protocols
@@
Fig. 5 shows the indoor experimental setup. Ground-truth lower-limb kinematics were captured using an optical motion capture system (10 Vicon Bonita cameras from Vicon Ltd.) A three-segment walkway ($7.5$ m long, $0.76$ m wide, and $0.14$ m high) was constructed for indoor experiments. Segment $1$ was built with reinforced plywood with embedded force plate #$\#1$ (model ACG-O from AMT Inc.) to replicate a solid ground condition. Segment $2$ was a sand-filled box with buried force plate #2 (from Bertec Corp.) 14 cm beneath the surface. GRF measurements were calibrated for sand layer thickness [6]. Segment $3$, made of the same material as Segment $1$, ensured natural gait transition from sand. The three-segment walkway design ensures that the subjects employ the consistent walking strategy throughout the trial. For evaluation purposes, subjects were equipped with eight wireless EMG sensors (from Delsys Inc.) on the right leg, monitoring rectus femoris (RF), vastus lateralis (VL), vastus medialis (VM), biceps femoris (BF), semitendinosus (SEM), tibialis anterior (TA), lateral gastrocnemius (LG), and medial gastrocnemius (MG). The EMG data were sampled at $1500$ Hz, followed by notch filtering ($58$-$62$ Hz) to remove line noise and band-pass filtering ($30$-$500$ Hz with $4$th-order zero-phase Butter-worth filters).
IV-B Indoor/Outdoor Tests and Data Collection
@@
Key evaluation metrics included: (1) knee joint torque estimation and GRF prediction accuracy; (2) impact on lower-limb muscle activation; and (3) controller performance in reducing metabolic costs on different terrains. Results from indoor experiments were used to address the first two objectives, while outdoor experiments were conducted to evaluate metabolic costs over extended periods. To obtain accurate subject-specific dynamic models in (3), a system identification is required for each subject. To do so, two experiment trials were conducted for each subject. The first trial identified the subjects’ dynamics without the exoskeleton, while the second experiment for the human-exoskeleton system. Detailed experiment procedures and the optimization method can be found in [12].
@@
V Results
V-A Gait Kinematics and Kinetics
The average self-selected walking speed across participants was approximately $1.02\pm 0.08$ m/s. Fig. 6 shows the comparison between the actual and predicted GRFs (normalized by body weight) on sand in indoor experiments. The prediction root mean square errors (RMSE) for normalized $F_{x}$ was $0.052\pm 0.015$ on solid ground and $0.12\pm 0.027$ on sand. For normalized $F_{z}$, the RMSE was $0.11\pm 0.018$ on solid ground and $0.11\pm 0.021$ on sand. The observed higher variance in GRF predictions on sand was attributed to the deformable terrain. The lowest RMSE was observed for the $F_{x}$ component on solid ground. To statistically evaluate the differences in GRFs across terrains, a paired sample t-test was conducted and the analysis revealed a statistically significant difference ($p<0.05$) in the GRFs on solid ground and sand. This finding supports the hypothesis that the GRFs can be used for terrain classifications and terrain variability significantly influences the accuracy of GRF predictions.
@@ V-A Gait Kinematics and Kinetics
Figs. 7 and 7 show the human walking locomotion parameters such as step length, width, and the center of mass (COM) variation of human body on two different terrains. The COM variation is defined as the subject’s normalized COM height deviation in the anterior-posterior direction during the gait cycle. It is clear to see an increase in the step length when human subjects wore the knee exoskeletons on the solid ground while the increase became less significant on sand. Paired sample t-tests were conducted to analyze the differences between each group. The proposed MPC enabled subjects to walk with the widest step width on sand and the narrowest on solid ground, while maintaining minimal COM variation on both terrains.
@@
Fig. 8 shows the average knee joint angle and torque (normalized by body weight) of four groups on the solid ground and sand. The use of exoskeleton aims at reducing human effort rather than replacing natural knee torque entirely. The overall knee kinematics show no significant deviation across the conditions, particularly in the early stance phase (i.e., 0-20% of the gait phase). Only a relatively flat knee flexion angle during stance phase (i.e., 0065656565% of gait phase) and a decrease in maximum flexion during swing phase (i.e., 65656565100100100100% of gait phase) on sand were observed. Regarding the exoskeleton torque, both the baseline controller and proposed MPC stiffness controller provided the more torque assistance during the stance phase than for the swing phase. Furthermore, the exoskeleton assisted human walking more significantly within the late stance phase (i.e., 3535353565656565% of gait phase) on sand than on solid ground. It is also found that the MPC controller offered larger joint moment assistance on sand than that by the baseline controller during the swing phase.
@@
Figure 8: Average knee angle, moment, and exoskeleton torque results over one complete gait cycle for solid ground (left column) and sand (right column).
Figure 10: Normalized average (a) RMS and (b) maximum EMG activation under four conditions for all subjects on solid ground and sand. Statistical analysis was conducted between Group D and the other groups on each terrain. The label * indicates a significant ($p<0.05$) difference between groups.
Figs. 9 and 9 show the EMG signals for eight muscles from a representative subject, normalized with respect to those of Group A on solid ground. Both RMS and max EMG metrics are included because RMS provides a measure of the overall muscle activity over time, while max EMG captures the peak activation. Peak EMG signals were highest in the unpowered condition (Group B) for both terrains, attributed to increased mass and resistance from the exoskeleton. On solid ground, both controllers showed comparable muscle activation reductions. However, the MPC led to notable decreases in peak EMG for ankle extensors and flexors (TA, LG, and MG), particularly on sand. Walking on granular terrain requires coordinated control across the lower limbs. Although torque was applied at the knee, the reduced muscle activation at the ankle shows that the exoskeleton assistance had a positive effect on the dynamics of the entire lower limb. This indicates improved gait stabilization and reduced physical effort across multiple joints, not just the knee.
@@
The RMS and maximum EMG values have been normalized to those of Group A on solid ground, accounting for all trials and subjects across the four groups. Figs. 10 and 10 show the results of this analysis. Among all trials, Group B consistently exhibited the highest activity on both terrains. Paired t-tests revealed statistically significant decreases in RF, VL, LG, and MG muscles for Group D on both terrains. Table I summarizes the RMS and maximum muscle activity percentage changes for Groups C and D relative to Group A, where positive values indicate an increase in muscle activity, while negative values represent a reduction. Both controllers demonstrated similar trends in alleviating muscle activation across terrains. On solid ground, Group C showed slight advantages in VL and BF RMS values, while Group D significantly reduced LG and MG activation. On sand terrain, Group D showed pronounced reductions, particularly in RF, TA, LG, and MG RMS values. Although the SEM activation increased on sand for both groups, Group D showed smaller increases in both RMS and maximum values, suggesting consistent performance and adaptive potential across terrains.
TABLE I: RMS/Max EMG comparison in percentage change (%)

### v2

@@ Abstract
Human walkers encounter diverse terrains, such as sand and solid ground, which result in varied gait locomotion and energy costs. This study aims to design and validate a stiffness-based model predictive control approach for knee exoskeleton assistance with walking on sand. A comparative analysis of gait and locomotion on sand versus solid ground is conducted. A machine learning-based estimation scheme is developed to predict ground reaction forces in real time. These predictions, combined with human joint torque estimates, are used to design a knee exoskeleton controller that employs a model predictive stiffness control strategy. The experiments demonstrate significant differences in lower limb kinematics and kinetics. Reductions in major muscle activation and metabolic cost during assisted walking on sand are also observed with $15\%$ and $3.7\%$, respectively. The proposed knee exoskeleton control framework improves walking efficiency on granular terrains by reducing both muscular and metabolic demands.
{IEEEkeywords}
@@
Human locomotion, granular terrain, exoskeleton control, model predictive control.
1 Introduction
@@
Human often traverse real-world environments with a variety of terrains that can disrupt steady gait and require additional effort [1]. While significant research has focused on walking on solid ground [2], real-world locomotion often occurs on granular terrains like sand, and this poses unique challenges due to the terrain’s deformable and shifting properties [3]. Granular terrains significantly alter the biomechanic characteristics of walking gait, including joint-level kinematics, kinetics, and energy expenditure. Studies have shown that walking on sand requires large ankle plantarflexion in stance phase and increased knee and hip flexion during leg swing [4]. Substantial changes on walking kinematics and greater postural instability were observed for walking on sand [5]. Similar studies of walking on compliant substrates demonstrated elevated activity and great mechanical work through large excursion and maximum flexion in hip and knee joints [6]. It has also been shown that walking on sand increases energy and metabolic cost [7, 6], shifts center-of-pressure trajectories [8], and induces greater mechanical work [9]. The work in [10] offered ground reaction forces and gait datasets with biomechanics of walking on solid ground and sand. Recent investigations into foot-substrate interactions highlight the influence of foot shape and speed on energy expenditure [11, 12]. These studies underscore the importance of terrain-specific insights to understand and optimize human walking locomotion.
@@
Although walking on sand has been explored, limited studies are reported for using wearable assistive devices such as exoskeletons for enhancing the walking stability and reducing metabolic expenditure. Recent advancements in wearable exoskeletons and prosthesis development have shown promising capability in restoring ambulatory function [13], aiding postural adjustment, reducing muscle load [14], and decreasing metabolic cost during walking locomotion [15]. Knee exoskeletons have emerged as effective tools for enhancing mobility and stability, particularly in individuals with gait impairments or during challenging locomotion tasks [16]. Recent studies show that knee exoskeleton-assisted walking improves joint kinematics and enhance gait efficiency and stability by modulating neural activation and reducing muscle effort in individuals with impaired motor control [17, 18]. Various control methods have been developed, including quasi-stiffness approaches that match target profiles [19] and reinforcement learning for impedance control [20]. Novel impedance modulation strategies have been proposed for lower-limb exoskeletons to assist sit-to-stand movements while preserving human control [21], and for soft exoskeletons to aid push-off on various terrains [22]. Trajectory tracking controls have been applied to lower limb exoskeletons to personalize and stabilize user gait movements with adaptive strategies [23, 24]. Model predictive control (MPC) was developed for compliant device behavior [25] and on-the-fly transitions between assistance modes [26, 27].
@@
All aforementioned exoskeleton applications were primarily designed for human walking on solid ground, and the developed controllers cannot be directly applied to granular terrains. Moreover, the effect of knee exoskeleton control on joint moments among the ankle, knee, and hip joints on sand remains largely unexplored. To bridge this gap, we hypothesize that: (1) knee robotic assistance can reduce muscle activation and energy expenditure during walking on sand; and (2) knee exoskeleton control would alter the distribution of joint moments and joint angles among the ankle, knee, and hip joints for adapted walking efficiency. To test these hypotheses, we focus on designing and evaluating a knee exoskeleton for walking assistance on sandy terrains, aiming to enhance ambulation ability while reducing metabolic cost. We propose a model predictive stiffness control design of human-exoskeletons to provide assistive knee joint torque and regulate walking gait on sand. A stiffness-based computation of desired knee torque is first designed using the observations of walking locomotion on sand and solid ground. A neural network is used to model human foot-terrain interactions and compute resultant ground reaction forces (GRFs) from measurements by wearable inertial measurement units (IMUs). Both desired knee torque and external GRFs are treated as inputs of the human-exoskeleton dynamic model. Finally, an MPC framework is formulated to regulate assistive torque throughout the gait phase. Both indoor terrain and outdoor experiments are conducted to validate the controller performance.
The contributions of this study are twofold: (1) a detailed analysis of joint-level biomechanics during walking on solid ground and sand, focusing on GRFs, joint torques, and energy cost; and (2) the design and validation of an adaptive exoskeleton control framework for granular terrains, emphasizing biomechanical insights and real-time adaptability. By addressing these biomechanical challenges, this study not only extends the applicability of wearable assistive devices to diverse environments but also provides novel insights into the dynamics of human-exoskeleton interaction on yielding terrains. The proposed stiffness-based MPC framework improves energy efficiency and gait stability on compliant surfaces, offering significant implications for assistive device design and rehabilitation. By examining the distribution of joint moments and angles across the hip, knee, and ankle joints, we highlight the role of knee assistance in redistributing mechanical effort and reducing metabolic cost.
2 Exoskeleton-Assisted Human Walking on Granular Terrain
2.1 Gait and Locomotion on Granular Terrains
Compared to solid ground, human walking on sand shows different gait and locomotion characteristics. Fig. 1 shows the human subject with wearable IMUs and bilateral knee exoskeletons that are used in experiments. Fig. 1 illustrates the kinematics and modeling configuration of the human-exoskeleton interactions on sand. A two-link limb model is used with joint angle $\boldsymbol{q}=[\theta_{t}\;\theta_{k}]^{T}\in\mathbb{R}^{2}$, where $\theta_{t}$ and $\theta_{k}$ represent the thigh and knee angle, respectively. The longitudinal and vertical GRFs are denoted as $F_{x}$ and $F_{z}$, respectively. The walking gait progression, namely, the joint angles and moments, was normalized by a gait phase variable $s$ (between 0 and 1) according to the heel strike events.
@@ 2.1 Gait and Locomotion on Granular Terrains
2.2 Learning-based GRF Estimation
Estimating GRFs is essential for real-time exoskeleton control on sand, as direct measurements are impractical. A machine learning-based model predicts GRFs using temporal and spatial features from IMU data, offering a non-invasive, real-time solution by capturing motion and force dynamics without ground contact. Fig. 3 illustrates a dual-pathway machine learning model for real-time GRF prediction. Input data are from IMUs mounted on the stance leg’s shank, heel, and toe; see Fig. 1. Similar to [28], only the accelerations in the sagittal plane ($x$- and $z$-axis) and the frontal-axis gyroscopic measurements ($y$-axis) from each IMU are used as the input for the machine learning model, i.e., $\boldsymbol{X}_{t}=[a_{x}(t)\;a_{z}(t)\;\omega_{y}(t)]^{T}$ from all three IMUs at time $t$.
@@ 2.2 Learning-based GRF Estimation
Outputs from LSTM and MLP are combined via dot product, emphasizing temporal-spatial feature interaction for a more accurate GRF estimation. A final dense layer with linear activation predicts GRFs. The model is trained using mean squared error loss and Adam optimizer, with early stopping and learning rate scheduling for regularization and optimization. This architecture is hypothesized to outperform traditional single-pathway models [29] by explicitly modeling both temporal progression and spatial configurations.
2.3 Stiffness-Based Exoskeleton Control
Fig. 4 illustrates the overall knee exoskeleton control design. The control scheme is built on real-time estimation of human torque $\hat{\boldsymbol{\tau}}_{h}$ from the stiffness-based model and external torque $\hat{\boldsymbol{\tau}}_{ext}$ via a machine learning model of the GRFs. IMUs and encoders provide the gait phase variable $s$ and lower-limb joint kinematics. An MPC scheme is formulated to design the controlled exoskeleton torque, denoted by $\boldsymbol{\tau}_{e}$. The lower-level controller of the exoskeletons takes the MPC outcome to drive the human-exoskeleton system to maintain desired joint angle and torque profiles on different terrains. The exoskeleton lower-level control is designed as a proportional-integral (PI) controller and the feedback torque signal is from the motor current sensor.
@@ 2.3 Stiffness-Based Exoskeleton Control
3 Experimental Design
3.1 Experiment Setup and Protocols
@@
Fig. 5 shows the indoor experimental setup. Ground-truth lower-limb kinematics were captured using an optical motion capture system (10 Vicon Bonita cameras from Vicon Ltd.) A three-segment walkway ($7.5$ m long, $0.76$ m wide, and $0.14$ m high) was constructed for indoor experiments. Segment $1$ was built with reinforced plywood with embedded force plate #$\#1$ (model ACG-O from AMT Inc.) to replicate a solid ground condition. Segment $2$ was a sand-filled box with buried force plate #2 (from Bertec Corp.) 14 cm beneath the surface. GRF measurements were calibrated for sand layer thickness [10]. Segment $3$ was made of the same material as Segment $1$ to allow natural gait transition from sand. The three-segment walkway design ensures that the subjects employ the consistent walking strategy throughout the trial. For evaluation purposes, subjects were equipped with eight wireless EMG sensors (from Delsys Inc.) on the right leg, monitoring rectus femoris (RF), vastus lateralis (VL), vastus medialis (VM), biceps femoris (BF), semitendinosus (SEM), tibialis anterior (TA), lateral gastrocnemius (LG), and medial gastrocnemius (MG). The EMG data were sampled at $1500$ Hz, followed by notch filtering ($58$-$62$ Hz) to remove line noise and band-pass filtering ($30$-$500$ Hz with $4$th-order zero-phase Butter-worth filters).
3.2 Indoor/Outdoor Tests and Data Collection
@@
Key evaluation metrics included: (1) GRF prediction accuracy; (2) controller’s influence on joint kinematics and kinetics; (3) impacts on lower-limb muscle activation; and (4) performance in reducing metabolic costs on different terrains. Results from indoor experiments were used to address the first three objectives, while outdoor experiments were conducted to evaluate metabolic costs over extended periods. To obtain accurate subject-specific dynamic models in (3), a system identification is required for each subject. To do so, two experiment trials were conducted for each subject. The first trial identified the subjects’ dynamics without the exoskeleton, while the second experiment for the human-exoskeleton system. Detailed experiment procedures and the optimization method can be found in [21].
@@
4 Results
4.1 Gait Kinematics and Kinetics
The average self-selected walking speed across participants was approximately $1.02\pm 0.08$ m/s. Fig. 6 shows the comparison between the actual and predicted GRFs (normalized by body weight) on sand in indoor experiments. The contact phase refers to the portion of the gait cycle during which the foot maintains contact with the ground. The prediction root mean square errors (RMSE) for normalized $F_{x}$ was $0.052\pm 0.015$ on solid ground and $0.12\pm 0.027$ on sand. For normalized $F_{z}$, the RMSE was $0.11\pm 0.018$ on solid ground and $0.11\pm 0.021$ on sand. The observed higher variance in GRF predictions on sand was attributed to the deformable terrain. The lowest RMSE was observed for the $F_{x}$ component on solid ground. To statistically evaluate the differences in GRFs across terrains, a paired sample t-test was conducted and the analysis revealed a statistically significant difference ($p<0.05$) in the GRFs on solid ground and sand. This finding supports that the GRFs can be used for terrain classifications and terrain variability significantly influences the accuracy of GRF predictions.
@@ 4.1 Gait Kinematics and Kinetics
Fig. 7 shows a clear comparison of average exoskeleton torque outputs between Groups C and D. The use of exoskeleton aims at reducing human effort rather than replacing natural knee torque entirely. Both the baseline controller and the MPC stiffness controller provided the more torque assistance during the stance phase than for the swing phase. Furthermore, the exoskeleton assisted human walking more significantly within the late stance phase (i.e., 3535353565656565% of gait phase) on sand than on solid ground. It is also found that the MPC controller offered larger joint moment assistance on sand than that by the baseline controller during the swing phase.
@@
Figure 7: Average exoskeleton torque outputs comparison between Groups C and D on solid ground (left) and sand (right).
Figure 9: (a) Comparison of hip, knee, and ankle joint angles over one complete gait cycle for solid ground (left column) and sand (right column). The top row: hip flexion(+)/extension(-); the mid row: flexion(+)/extension(-); the bottom row: ankle dorsiflexion(+)/plantarflexion(-). The thick curves are the mean value profiles, while the shaded ares show the one standard deviation around the mean values. (b) The results of 1DSPM statistical analyses comparing walking with the baseline controller (left) vs with the MPC controller (right) on sand. The black curves represent the test statistic (t-value) across the gait cycle, and the red dashed lines indicate the critical threshold for statistical significance at $\alpha=0.05$. Shaded regions above the threshold indicate phases of the gait cycle where significant differences were detected. Stars (*, ,**,∗ ∗ , )***)∗ ∗ ∗ ) denote p-values of $<$0.05, $<$0.01, and $<$0.001, respectively. The x-axis represents the gait phase, normalized from 0 (heel strike) to 1 (subsequent heel strike) for a complete gait cycle.
Figure 10: Comparison of hip, knee, and ankle joint moments over one complete gait cycle for solid ground (left column) and sand (right column). The top row: hip flexion(+)/extension(-); the mid row: flexion(+)/extension(-); the bottom row: ankle dorsiflexion(+)/plantarflexion(-). The thick curves are the mean value profiles, while the shaded ares show the one standard deviation around the mean values. (b) The results of 1DSPM statistical analyses comparing walking with the baseline controller (left) vs with the MPC controller (right) on sand. The black curves represent the test statistic (t-value) across the gait cycle, and the red dashed lines indicate the critical threshold for statistical significance at $\alpha=0.05$. Shaded regions above the threshold indicate phases of the gait cycle where significant differences were detected. Stars (*, ,**,∗ ∗ , )***)∗ ∗ ∗ ) denote p-values of $<$0.05, $<$0.01, and $<$0.001, respectively. The x-axis represents the gait phase, normalized from 0 (heel strike) to 1 (subsequent heel strike) for a complete gait cycle.
Figs. 8 and 8 show the human walking locomotion parameters such as step length, width, and the center of mass (COM) variation of human body on two different terrains. The COM variation is defined as the subject’s normalized COM height deviation in the anterior-posterior direction during the gait cycle. It is clear to see an increase in the step length when human subjects wore the knee exoskeletons on the solid ground while the increase became less significant on sand. Paired sample t-tests were conducted to analyze the differences between each group. The MPC-based exoskeletons enabled subjects to walk with the widest step width on sand and the narrowest on solid ground, while maintaining minimal COM variation on both terrains.
Fig. 9 shows the hip, knee, and ankle joint angle profiles of four groups on the solid ground and sand. The hip joint angles exhibit greater flexion angles during the swing phase with the exosekelton. The knee kinematics show no significant deviation in the early stance phase (i.e., 0-20% of the gait phase). A relatively flat knee flexion angle during stance phase (i.e., 0065656565% of gait phase) and a decrease in maximum flexion during swing phase (i.e., 65656565100100100100% of gait phase) on sand were observed. The ankle joint demonstrated greater plantarflexion during mid-stance on sand, particularly in Group C. We also conducted a statistical analysis across the entire gait cycle using one-dimensional statistical parametric mapping (1DSPM) to capture joint angle variations. The results comparing walking with the baseline controller (left) vs with the MPC controller (right) on sand are presented in Fig. 9. For the hip joint, two main regions of statistical significance ($p<0.001$) were observed within the gait cycle under both controller, i.e., during early to mid swing phase. Under the MPC controller, the knee joint exhibited two significant regions of statistical significance during the mid stance ($p<0.01$) and terminal swing phase ($p<0.05$) of the gait cycle. Regarding the ankle joint, no region of statistical significance throughout the gait is obtained under either controller.
Fig. 10 shows the joint torque profile (normalized by body weight) of the hip, knee, and ankle joints across the gait cycle on solid ground and sand. The 1DSPM analysis results between walking on sand with the baseline controller vs with MPC controller are also presented in Fig. 10. Although the exoskeleton applied the torques only at the knee joint, the joint torque profiles change significantly at the hip and ankle joints as well. The hip joint displayed an increased extension moment on both terrains with the exoskeleton during early swing phase, with walking on sand showing more significant increases and longer periods, corresponding to the larger hip flexion angles. On solid ground, the knee joint exhibited a characteristic biphasic pattern with peak extension moments during early stance and late stance. On sand, the MPC controller leads to higher knee joint moments during the late stance to early swing phase compared to both the unpowered exoskeleton (Group B) and the baseline controller (Group C). The knee moments during late stance and early swing phase was significantly increased across all conditions ($p<$ 0.001). In contrast, the ankle joint moments demonstrated increased variability on sand, particularly during push-off. Group D showed a significant reduction ($p<$ 0.001) in peak ankle plantarflexion moments compared to other three groups, while the baseline controller increased the plantarflexion moments significantly ($p<$ 0.001).
@@
Figure 12: Normalized average (a) RMS and (b) maximum EMG activation under four conditions for all subjects on both terrains. Statistical analysis was conducted between Group D and the other groups on each terrain. The label * indicates a significant ($p<0.05$) difference between groups.
4.2 Muscle Activation and Metabolic Cost
Figs. 11 and 11 show the EMG signals for eight muscles from a representative subject, normalized with respect to those of Group A on solid ground. Both RMS and maximum EMG metrics are included because RMS provides a measure of the overall muscle activity over time, while maximum EMG captures the peak activation. Peak EMG signals were highest in the unpowered condition (Group B) for both terrains, attributed to increased mass and resistance from the exoskeleton. On solid ground, both controllers showed comparable muscle activation reductions. However, the MPC led to notable decreases in peak EMG for ankle extensors and flexors (TA, LG, and MG), particularly on sand. Walking on granular terrain requires coordinated control across the lower limbs. Although torque was applied at the knee, the reduced muscle activation at the ankle shows that the exoskeleton assistance had a positive effect on the dynamics of the entire lower limb.
The RMS and maximum EMG values have been normalized to those of Group A on solid ground, accounting for all trials and subjects across the four groups. Figs. 12 and 12 show the results of this analysis. Among all trials, Group B consistently exhibited the highest activity on both terrains. Statistical analysis revealed significant decreases in RF, VL, LG, and MG muscles for Group D on both terrains. Table 1 summarizes the RMS and maximum muscle activity percentage changes for Groups C and D relative to Group A, where positive values indicate an increase in muscle activity and negative values represent a reduction. Both controllers demonstrated similar trends in alleviating muscle activation across terrains. On solid ground, Group C showed slight advantages in VL and BF RMS values, while Group D significantly reduced LG and MG activation. On sand terrain, Group D showed pronounced reductions, particularly in RF, TA, LG, and MG RMS values. Although the SEM activation increased on sand for both groups, Group D showed smaller increases in both RMS and maximum values.
Table 1: RMS/Max EMG comparison in percentage change (%)